OpenStudy (anonymous):

Find the 110th term of the sequence -7, 3, 13, 23, ... Please Help! Medal and Fan!

OpenStudy (anonymous):

@dan815 @Hero @pooja195 Can any of y'all help please..?

OpenStudy (decarr432):

is this all thats given

OpenStudy (decarr432):

@Caseyyyy15

OpenStudy (anonymous):

Yeah thats it! @Decarr432

OpenStudy (jhannybean):

\[a_n = a_1 +(n-1)d\]\[a_{110}= -7+(110-1)d\]\[d: 3-(-7) =10 ~,~ 13-3 = 10~,~ 23-13 = 10~,~ ... ~\therefore ~d=10\]

OpenStudy (jhannybean):

\[a_{110} = -7 +(110-1)(10)\]

OpenStudy (anonymous):

Thank you so much @Jhannybean :)

OpenStudy (jhannybean):

Np